# Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images

### v2

@@ 3.2 View generation
Both backbones use a text-to-image stable diffusion model [RBL22] to generate candidate RGB images based on the input grids. The stable diffusion model gradually denoises a random normal noise image in latent space $\mathbf{z}\in^{h\times w\times l}$, where ${h=w=64}$ and $l=4$ are the stable diffusion’s latent space dimensions. The outputs of the diffusion model blocks are modulated by a ControlNet network branch [ZRA23], which is conditioned on the encoded text, depth map grid, and, depending on the specific backbone and iteration, on the rendered maps derived from partially textured shapes. The denoised latent is decoded into a grid of images $\{\mathbf{I}_{c}\}_{c=1}^{C}$ for the selected viewpoints:
@@ 3.3.1 Traditional backprojection
Previous methods use an inverse UV mapping procedure for backprojection. Specifically, given each texel in the texture map $\mathbf{u}=(u,v)\in\mathbf{T}$, its corresponding 3D surface point $\mathbf{s}_{\mathbf{u}}=(x_{\mathbf{u}},y_{\mathbf{u}},z_{\mathbf{u}})\in\mathbf{S}$ is first estimated. Practically, this can be implemented by rendering a flattened version of the input polygon mesh $\mathbf{S}$ with its vertex coordinates replaced with its texture coordinates. Then for each rendered pixel, its barycentric coordinates are calculated within the flattened triangle it belongs to. These are used to interpolate the 3D vertex positions of this triangle in the original mesh to acquire the corresponding 3D point $\mathbf{s}_{\mathbf{u}}$ for that texel. The procedure assumes that each texture coordinate maps to a single 3D face – if a texture coordinate is re-used by multiple faces, the texture can be unwrapped to avoid this [LPRM02].
@@
Most previous methods, such as Paint3D [ZCQ24], Text2Tex [CSL23], TEXTure [RMA23], find the view where the 3D point appears to be the most front-facing i.e., the dot product between its normal $\mathbf{n}_{\mathbf{u}}$ and the view vector $\mathbf{v}_{c}$ is maximized, and simply copy the color from the generated image pixel where the 3D point is projected onto under that view:
@@
where $\mathcal{R}_{c^{\prime}}$ returns the 2D pixel coordinates of the point $\mathbf{s}_{\mathbf{u}}$ rendered onto the image $\mathbf{I}_{c^{\prime}}$ under the most front-facing viewpoint $c^{\prime}$ for this point. It is also common to employ a hand-tuned threshold $\mathbf{n}_{\mathbf{u}}\cdot\mathbf{v}_{c}>thr$ to avoid copying colors from obscure views. We note some texels may not acquire any color, if their corresponding points are not accessible by any acceptable views – texture inpainting is used to fill such texels with color [ZCQ24]. Unfortunately, this strategy can easily lead to inconsistencies e.g., texels of neighboring 3D points might acquire colors from different views that may not blend well together.
@@
An alternative strategy, followed by MatAtlas [CDG24] in its first diffusion iteration, is to average colors from the pixels of all views accessing the texel’s corresponding point to blends any small inconsistencies:
@@
Other approaches [ZPZ24, CMZ24] implement a weighted averaging scheme, where the weights are the dot product between the 3D point normals $\mathbf{n}_{\mathbf{u}}$ and view vectors $\mathbf{v}_{c}$. Unfortunately, averaging schemes can yield blurry texture results, as also noted in [CDG24].
@@ 3.3.2 Neural backprojection
Instead of relying on ad hoc, hand-tuned schemes for backprojecting and blending colors from the generated views, we instead propose a learned backprojection scheme. We utilize a neural module based on attention [VSP17] to assign appropriate colors to each texel by comparing its features with those of pixels gathered from image neighborhoods related to this texel across all views. The texel and pixels features are learned based on positional encodings of the underlying 3D points corresponding to these texels and pixels as well as their underlying appearance (color). The positional encodings incorporate information about their 3D position, normals, angles between normals and view vectors, and surface coordinates encoded in geodesic distances – the reason for using all this information is that the texel color should not be determined by a pixel from a single view, or by merely averaging pixels, but instead by considering broader pixel neighborhoods across all views to maximize view consistency, and by considering texture correlations in local surface neighborhoods according to the underlying 3D geometry to promote texture consistency.
@@
For each pixel $\mathbf{p}\in\mathcal{N}(\mathbf{s}_{u})$, we determine the corresponding 3D surface point projected onto this pixel based on the view the pixel originated from. We then compute a feature vector that encodes the 3D position $\mathbf{s}_{\mathbf{p}}$ and normal $\mathbf{n}_{\mathbf{p}}$ of this surface point relative to the texel’s corresponding surface point. Pixels whose 3D locations are closer to the texel’s point, or have more similar normals, are expected to have a stronger influence on its color. Additionally, we encode the geodesic distance $\delta_{\mathbf{p},\mathbf{u}}$ between the pixel’s surface point and the texel’s 3D point. Geodesic distances refine pixel contributions by accounting for true surface proximity unlike Euclidean distances, which may misleadingly suggest closeness e.g., in regions with folds and high-curvature regions (Figure 7). Geodesic distances are computed using the method in [MR12]. The encoding is obtained via a trained MLP using the following features:
@@
Figure 3: Comparisons between our method, Im2SurfTex, and Paint3D [ZCQ24]. Paint3D suffers from view projection artifacts when there are steep depth changes or occluded regions in the input views, as its heuristic best view selection strategy leads to texture discontinuities and inconsistencies. In contrast, our approach generates more seamless and coherent textures.
@@
where $\mathbf{Q},\mathbf{K},\mathbf{V}$ are learned transformations. Note that the positional encodings are added to the rest of the features, as also done in [VSP17]. Note that in our case, the value transformation involves only the color encodings, as our end goal is to transform pixel colors (rather than position) to texel colors. Based on the query and key transformations, we compute the attention weights, which represent the importance of each pixel in contributing to the texel’s color:
@@ 3.4 Texture inpainting and refinement
After backprojection and the final iteration of either backbone, some texels may still remain empty. For texture inpainting, we follow Paint3D’s approach: a trained diffusion model fills any texture holes within the UV plane. Additionally, Paint3D’s high-definition (HD) diffusion model is subsequently used to further enhances the visual quality of the texture map in UV space. We refer readers to Paint3D [ZCQ24] for more details, and the authors’ implementation for these trained modules. We note that we apply the same inpainting and HD processing for both backbone implementations. Unfortunately, as shown in our experiments, these post-processing modules often fail to correct the artifacts introduced by traditional backprojection.
@@ 3.5 Training
We train the parameters of our MLPs and cross-attention module based on supervision from Objaverse [DSS23]. We use the training split from Paint3D, a subset of the Objaverse dataset containing approximately 100K shapes, each paired with a target texture image. We preprocess the data by computing geodesic neighborhoods for each object, storing the resulting tensors as additional shape-specific information. The network renders input views, which are then used to reconstruct the target texture during training. Our network is trained with a batch size of $4$ for $10$ epochs on four NVIDIA A6000 GPUs, taking approximately five days. To make our model more robust to any view inconsistencies, we employ a mixed batch approach where some renders are re-generated using a pretrained Stable Diffusion 1.5 model [RBL22] with partial noise levels ranging from $0.2$ to $0.7$. Samples with $0.2$ noise introduce minor variations, while those with $0.7$ noise introduce significant deviations from the target texture. For training, we optimize the model’s weights using an L1 loss function between the generated and target texture images.
@@ 3.6 Implementation details
During inference, our approach follows either backbone described in Section 3.1, yet incorporating the learned backprojection module instead of their heuristic backprojection. Since some texels remain unfilled after backprojection, they are subsequently inpainted and refined using pretrained Stable Diffusion 1.5 and the Paint3D’s 3D-aware ControlNet module. The final output textures have a resolution of $1024\times 1024$. The entire texturing process takes one to two minutes on a single NVIDIA A6000 GPU to texture an input shape. Each iteration of view generation and neural backprojection takes approximately $35$ seconds, while the inpainting and high-definition (HD) modules require around $20$ seconds each. For preprocessing, our approach employs a one-time procedure to compute geodesic information metadata, which takes approximately $30$ minutes when processing a new object for the first time. This part can be significantly accelerated with more efficient techniques for computation of geodesics [CWW13, ZHA23]. We also refer readers to our project page with source code for more details.  [^1] Project page (with code): ygeorg01.github.io/Im2SurfTex
Figure 4: Comparisons between our method and MatAtlas [CDG24]. MatAtlas struggles with inconsistencies in the output texture, particularly in regions with high curvature, where misalignments become more apparent. In contrast, as shown in the figure,  Im2SurfTex  tends to produce more coherent textures.
@@ 4.1 Experimental setup
We evaluate  Im2SurfTex on the test split provided by Text2Tex [CSL23]. The split includes $410$ textured meshes from Objaverse [DSS23] across $225$ categories. All competing methods are trained on the same training split, as discussed in Section 3.5, and evaluated on the same above test Objaverse split. We also note that all competing methods use the same UV maps and surface parametrization.
@@
For quantitative evaluation, we use standard image quality metrics for generative image models. Specifically, we report the Fréchet Inception Distance (FID[HRU17] and Kernel Inception Distance (KID[BSAG18]. The FID compares the mean and standard deviation of the deepest layer features in the Inception v3 network between the set of real and generated images. The KID calculates the maximum mean discrepancy between the real and generated images. In practice, the MMD is calculated over a number of subsets to obtain a mean and standard deviation measurement. Additionally, we measure alignment, or similarity, of the generated images with the input text prompt using the CLIP score [RKH21]. To compute these metrics, following [ZCQ24], we render each mesh with the generated textures from $20$ fixed viewpoints at a resolution of $512\times 512$. The reference distribution consists of renders of the same meshes using the textures found in the Objaverse dataset, under identical lighting settings.
@@ 4.2 Comparisons
In Table 2, we include quantitative comparisons of the best variant of our method (based on the MatAtlas backbone) with other state-of-the-art models for text-to-texture generation. Here we also include a comparison with the recent method of TEXGen[YYG24]. According to all the evaluation metrics, our method provides the best performance in terms of FID & KID distances as well as CLIP score.
In Figure 5, we show qualiitative comparisons with TEXGen’s released implementation [YYG24]. We observe that TEXGen often leads to global texture inconsistencies on the output shapes, while our method is more view-consistent.
@@
Figure 5: Comparison between  Im2SurfTex and TEXGen [YYG24]. We show two different views of the textured objects. Our method produces more coherent and view-consistent textures.
@@ 4.3 Ablation
Figure 8: Results for Paint3D and our method for $4$, $6$, and $8$ input views.  Im2SurfTex generates smoother surfaces.

### v3

@@ 3.2 View generation
Both backbones use a text-to-image stable diffusion model [Rombach:2022:High] to generate candidate RGB images based on the input grids. The stable diffusion model gradually denoises a random normal noise image in latent space $\mathbf{z}\in^{h\times w\times l}$, where h=w=64{h=w=64} and $l=4$ are the stable diffusion’s latent space dimensions. The outputs of the diffusion model blocks are modulated by a ControlNet network branch [Zhang:2023:ControlNet], which is conditioned on the encoded text, depth map grid, and, depending on the specific backbone and iteration, on the rendered maps derived from partially textured shapes. The denoised latent is decoded into a grid of images $\{\mathbf{I}_{c}\}_{c=1}^{C}$ for the selected viewpoints:
@@ 3.3.1 Traditional backprojection
Previous methods use an inverse UV mapping procedure for backprojection. Specifically, given each texel in the texture map $\mathbf{u}=(u,v)\in\mathbf{T}$, its corresponding 3D surface point $\mathbf{s}_{\mathbf{u}}=(x_{\mathbf{u}},y_{\mathbf{u}},z_{\mathbf{u}})\in\mathbf{S}$ is first estimated. Practically, this can be implemented by rendering a flattened version of the input polygon mesh $\mathbf{S}$ with its vertex coordinates replaced with its texture coordinates. Then for each rendered pixel, its barycentric coordinates are calculated within the flattened triangle it belongs to. These are used to interpolate the 3D vertex positions of this triangle in the original mesh to acquire the corresponding 3D point $\mathbf{s}_{\mathbf{u}}$ for that texel. The procedure assumes that each texture coordinate maps to a single 3D face – if a texture coordinate is re-used by multiple faces, the texture can be unwrapped to avoid this [Levy:2002:LSCM].
@@
Most previous methods, such as Paint3D [Zeng:2024:Paint3D], Text2Tex [Chen:2023:Text2tex], TEXTure [Richardson:2023:Texture], find the view where the 3D point appears to be the most front-facing i.e., the dot product between its normal $\mathbf{n}_{\mathbf{u}}$ and the view vector $\mathbf{v}_{c}$ is maximized, and simply copy the color from the generated image pixel where the 3D point is projected onto under that view:
@@
where $\mathcal{R}_{c^{\prime}}$ returns the 2D pixel coordinates of the point $\mathbf{s}_{\mathbf{u}}$ rendered onto the image $\mathbf{I}_{c^{\prime}}$ under the most front-facing viewpoint $c^{\prime}$ for this point. It is also common to employ a hand-tuned threshold $\mathbf{n}_{\mathbf{u}}\cdot\mathbf{v}_{c}>thr$ to avoid copying colors from obscure views. We note some texels may not acquire any color, if their corresponding points are not accessible by any acceptable views – texture inpainting is used to fill such texels with color [Zeng:2024:Paint3D]. Unfortunately, this strategy can easily lead to inconsistencies e.g., texels of neighboring 3D points might acquire colors from different views that may not blend well together.
@@
An alternative strategy, followed by MatAtlas [Ceylan:2024:Matatlas] in its first diffusion iteration, is to average colors from the pixels of all views accessing the texel’s corresponding point to blends any small inconsistencies:
@@
Other approaches [zhang2024texpainter, Cheng:2024:MVPaint] implement a weighted averaging scheme, where the weights are the dot product between the 3D point normals $\mathbf{n}_{\mathbf{u}}$ and view vectors $\mathbf{v}_{c}$. Unfortunately, averaging schemes can yield blurry texture results, as also noted in [Ceylan:2024:Matatlas].
@@ 3.3.2 Neural backprojection
Instead of relying on ad hoc, hand-tuned schemes for backprojecting and blending colors from the generated views, we instead propose a learned backprojection scheme. We utilize a neural module based on attention [Vaswani:2017:Attenion] to assign appropriate colors to each texel by comparing its features with those of pixels gathered from image neighborhoods related to this texel across all views. The texel and pixels features are learned based on positional encodings of the underlying 3D points corresponding to these texels and pixels as well as their underlying appearance (color). The positional encodings incorporate information about their 3D position, normals, angles between normals and view vectors, and surface coordinates encoded in geodesic distances – the reason for using all this information is that the texel color should not be determined by a pixel from a single view, or by merely averaging pixels, but instead by considering broader pixel neighborhoods across all views to maximize view consistency, and by considering texture correlations in local surface neighborhoods according to the underlying 3D geometry to promote texture consistency.
@@
For each pixel $\mathbf{p}\in\mathcal{N}(\mathbf{s}_{u})$, we determine the corresponding 3D surface point projected onto this pixel based on the view the pixel originated from. We then compute a feature vector that encodes the 3D position $\mathbf{s}_{\mathbf{p}}$ and normal $\mathbf{n}_{\mathbf{p}}$ of this surface point relative to the texel’s corresponding surface point. Pixels whose 3D locations are closer to the texel’s point, or have more similar normals, are expected to have a stronger influence on its color. Additionally, we encode the geodesic distance $\delta_{\mathbf{p},\mathbf{u}}$ between the pixel’s surface point and the texel’s 3D point. Geodesic distances refine pixel contributions by accounting for true surface proximity unlike Euclidean distances, which may misleadingly suggest closeness e.g., in regions with folds and high-curvature regions (Figure 7). Geodesic distances are computed using the method in [Melvaer:2012:GPC]. The encoding is obtained via a trained MLP using the following features:
@@
Figure 3: Comparisons between our method, Im2SurfTex, and Paint3D [Zeng:2024:Paint3D]. Paint3D suffers from view projection artifacts when there are steep depth changes or occluded regions in the input views, as its heuristic best view selection strategy leads to texture discontinuities and inconsistencies. In contrast, our approach generates more seamless and coherent textures.
@@
where $\mathbf{Q},\mathbf{K},\mathbf{V}$ are learned transformations. Note that the positional encodings are added to the rest of the features, as also done in [Vaswani:2017:Attenion]. Note that in our case, the value transformation involves only the color encodings, as our end goal is to transform pixel colors (rather than position) to texel colors. Based on the query and key transformations, we compute the attention weights, which represent the importance of each pixel in contributing to the texel’s color:
@@ 3.4 Texture inpainting and refinement
After backprojection and the final iteration of either backbone, some texels may still remain empty. For texture inpainting, we follow Paint3D’s approach: a trained diffusion model fills any texture holes within the UV plane. Additionally, Paint3D’s high-definition (HD) diffusion model is subsequently used to further enhances the visual quality of the texture map in UV space. We refer readers to Paint3D [Zeng:2024:Paint3D] for more details, and the authors’ implementation for these trained modules. We note that we apply the same inpainting and HD processing for both backbone implementations. Unfortunately, as shown in our experiments, these post-processing modules often fail to correct the artifacts introduced by traditional backprojection.
@@ 3.5 Training
We train the parameters of our MLPs and cross-attention module based on supervision from Objaverse [Deitke:2023:Objaverse]. We use the training split from Paint3D, a subset of the Objaverse dataset containing approximately 100K shapes, each paired with a target texture image. We preprocess the data by computing geodesic neighborhoods for each object, storing the resulting tensors as additional shape-specific information. The network renders input views, which are then used to reconstruct the target texture during training. Our network is trained with a batch size of $4$ for $10$ epochs on four NVIDIA A6000 GPUs, taking approximately five days. To make our model more robust to any view inconsistencies, we employ a mixed batch approach where some renders are re-generated using a pretrained Stable Diffusion 1.5 model [Rombach:2022:High] with partial noise levels ranging from $0.2$ to $0.7$. Samples with $0.2$ noise introduce minor variations, while those with $0.7$ noise introduce significant deviations from the target texture. For training, we optimize the model’s weights using an L1 loss function between the generated and target texture images.
@@ 3.6 Implementation details
During inference, our approach follows either backbone described in Section 3.1, yet incorporating the learned backprojection module instead of their heuristic backprojection. Since some texels remain unfilled after backprojection, they are subsequently inpainted and refined using pretrained Stable Diffusion 1.5 and the Paint3D’s 3D-aware ControlNet module. The final output textures have a resolution of $1024\times 1024$. The entire texturing process takes one to two minutes on a single NVIDIA A6000 GPU to texture an input shape. Each iteration of view generation and neural backprojection takes approximately $35$ seconds, while the inpainting and high-definition (HD) modules require around $20$ seconds each. For preprocessing, our approach employs a one-time procedure to compute geodesic information metadata, which takes approximately $30$ minutes when processing a new object for the first time. This part can be significantly accelerated with more efficient techniques for computation of geodesics [Crane:2013:Heat, Zhang:2023:NeuroGF]. We also refer readers to our project page with source code for more details.  [^1] Project page (with code): ygeorg01.github.io/Im2SurfTex
Figure 4: Comparisons between our method and MatAtlas [Ceylan:2024:Matatlas]. MatAtlas struggles with inconsistencies in the output texture, particularly in regions with high curvature, where misalignments become more apparent. In contrast, as shown in the figure,  Im2SurfTex  tends to produce more coherent textures.
@@ 4.1 Experimental setup
We evaluate  Im2SurfTex on the test split provided by Text2Tex [Chen:2023:Text2tex]. The split includes $410$ textured meshes from Objaverse [Deitke:2023:Objaverse] across $225$ categories. All competing methods are trained on the same training split, as discussed in Section 3.5, and evaluated on the same above test Objaverse split. We also note that all competing methods use the same UV maps and surface parametrization.
@@
For quantitative evaluation, we use standard image quality metrics for generative image models. Specifically, we report the Fréchet Inception Distance (FID[Heusel:2017:GANs] and Kernel Inception Distance (KID[Bińkowski:2018d:Demystifying]. The FID compares the mean and standard deviation of the deepest layer features in the Inception v3 network between the set of real and generated images. The KID calculates the maximum mean discrepancy between the real and generated images. In practice, the MMD is calculated over a number of subsets to obtain a mean and standard deviation measurement. Additionally, we measure alignment, or similarity, of the generated images with the input text prompt using the CLIP score [Radford:2021:CLIP]. To compute these metrics, following [Zeng:2024:Paint3D], we render each mesh with the generated textures from $20$ fixed viewpoints at a resolution of $512\times 512$. The reference distribution consists of renders of the same meshes using the textures found in the Objaverse dataset, under identical lighting settings.
@@ 4.2 Comparisons
In Table 2, we include quantitative comparisons of the best variant of our method (based on the MatAtlas backbone) with other state-of-the-art models for text-to-texture generation. Here we also include a comparison with the recent method of TEXGen[Yu:2024:TEXGen]. According to all the evaluation metrics, our method provides the best performance in terms of FID & KID distances as well as CLIP score.
In Figure 5, we show qualiitative comparisons with TEXGen’s released implementation [Yu:2024:TEXGen]. We observe that TEXGen often leads to global texture inconsistencies on the output shapes, while our method is more view-consistent.
@@
Figure 5: Comparison between  Im2SurfTex and TEXGen [Yu:2024:TEXGen]. We show two different views of the textured objects. Our method produces more coherent and view-consistent textures.
@@ 4.3 Ablation
Figure 8: Results for Paint3D and our method for 44, 66, and 88 input views.  Im2SurfTex generates smoother surfaces.